Find the next 3 terms in the sequence: 1, 4, 9, 16, ...
Thanks!

Answer by FrankM(1040)


1*1 = 1
2*2 = 4
3*3 = 9
4*4 = 16
5*5 = 25
6*6 = 36
7*7 = 49
25, 36, 49